OpenStudy (anonymous):

how many intersection do the equations x^2+y^2=9 and 3x^2+4y^2=16 have?

OpenStudy (lgbasallote):

use systems of equations..does that help?

OpenStudy (anonymous):

Kind of does

OpenStudy (lgbasallote):

wonderful

OpenStudy (anonymous):

Yes and i got 3 is that correct

OpenStudy (anonymous):

i do no think they even intersects !

OpenStudy (anonymous):

So i must have done something wrong but how can you tell

OpenStudy (anonymous):

the first equation is that of circle (centered at origin) of radius 3. the second equation is ellipse (centered at origin) and who major axis (4/sqrt(3)) is less than the radius of the circle (3) . so it is ellipse inside the circle !!!

OpenStudy (anonymous):

which many none intersection

OpenStudy (anonymous):

mean

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok. Thanks

OpenStudy (anonymous):

here they are .

OpenStudy (anonymous):

oK. i See i guess it will be better to draw it out

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Thanks

OpenStudy (lgbasallote):

@sami-21 you sure about that?

OpenStudy (anonymous):

100 % :P

OpenStudy (anonymous):

@lgbasallote there another way solve equation 1 for x^2 x^2=9-y^2 second equation becomes 3(9-y^2)+4y^2=16 y^2=-16 and you do not get real solutions !!! so there is no intersections a all .